Staff.

LBTO: Gustavo and Juan Carlos

summary.

Check of the new fiber for iLocator. The fiber can not be connected and it put it back the default one.

Measurements taken in open loop. System check before and after.

Details.

12:36 Power on WFS
  • OK.
  • Check ocam bias.
  • _193638
  • Bias OK.
  • cool down ocam.
  • dark ocam
20:03:35 PRESETAO. MAG 11

taking measurements.
  • 93phtyo/subap, mag 10.4
  • _211138
  • 10l/D
  • 78.1 phot/subap, mag 10.6
  • _211305
2:32pm switch fibers.

Gustavo realized that the connector is not the correct one. Fiber put it back to default.

Check afte the reinstallation.

Argos Cal Unit shutter full open

PresetAO mag 11

  • center.
  • dark
  • _233229,_233320
  • check flux.
  • 90 pho/subap mag 10.4
  • 10l/D
  • 83phot/subap mag 10.5
  • _233501,_233511

-- JuanCarlosGuerra - 08 Jul 2021
Topic revision: r4 - 09 Jul 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback